# M3Bench: Benchmarking Whole-body Motion Generation for Mobile Manipulation in 3D Scenes

### v2

@@ Abstract
We propose M3Bench, a new benchmark of whole-body motion generation for mobile manipulation tasks. Given a 3D scene context, M3Bench requires an embodied agent to understand its configuration, environmental constraints and task objectives, then generate coordinated whole-body motion trajectories for object rearrangement tasks. M3Bench features 30k object rearrangement tasks across 119 diverse scenes, providing expert demonstrations generated by our newly developed M3BenchMaker. This automatic data generation tool produces coordinated whole-body motion trajectories from high-level task instructions, requiring only basic scene and robot information. Our benchmark incorporates various task splits to assess generalization across different dimensions and leverages realistic physics simulation for trajectory evaluation. Through extensive experimental analyses, we reveal that state-of-the-art models still struggle with coordinated base-arm motion while adhering to environment-context and task-specific constraints, highlighting the need to develop new models that address this gap. Through M3Bench, we aim to facilitate future robotics research towards more adaptive and capable mobile manipulation in diverse, real-world environments.
@@ I Introduction
Humans possess an innate ability to manipulate their environment with remarkable flexibility and coordination, seamlessly integrating locomotion and manipulation. In contrast, robots still struggle to match this level of adaptability and proficiency in mobile manipulation tasks. Traditional trajectory optimization methods[1, 2], while capable of producing complex whole-body motion, are limited by their reliance on perfect environmental knowledge and predefined goal configurations (e.g., grasp poses), restricting their effectiveness and generalizability in real-world scenarios.
To address these limitations, researchers have turned to learning-based approaches. However, existing learning-based models deployed on mobile manipulators often address individual subproblems such as navigation, grasping, or motion planning in a modular fashion [3, 4]. These approaches can lead to suboptimal or infeasible solutions as they overlook the potential of coordinated whole-body motion and create misalignments between module outputs and requirements. For instance, in an object fetching task, a nearby navigable position might be impractical for the arm to reach the object, or a feasible grasp pose could be unachievable due to embodiment constraints (see Fig. 2(a)). These limitations underscore the need for a holistic approach that jointly understands robot embodiment, environment, and task objectives to generate feasible coordinated whole-body motion for mobile manipulators in diverse 3D scenes.
However, the development of such a holistic approach has been hindered by the scarcity of datasets capturing whole-body motions in 3D scenes. Existing benchmarks primarily focus on navigation [5] or task planning without actual object manipulation [6], lacking necessary data for studying embodied motion in 3D environments. Manipulation-oriented benchmarks often concentrate on object-centric modeling (e.g., grasping [7] and affordance [8, 9, 10]) with simplified embodiment, or stationary manipulation [11, 12] that neglects the coordination between mobile base and arm motions in diverse scenes.
To address this need, we developed M3BenchMaker (see Fig. 2(b)), an automatic data generation tool to produce whole-body motion trajectories as demonstrations for robot learning. M3BenchMaker procedurally generates coordinated trajectories from high-level task instructions, requiring only the action type, object link, and Unified Robot Description Format (URDF) of the scene and robot. It employs an energy-based model [13] to predict grasp pose or placement candidates and leverages an advanced virtual kinematics technique [1] to solve for coordinated whole-body motion trajectories (see Sec. III for details). This tool not only streamlines the demonstration collection process but also allows researchers to generate additional samples from customized robot and scene configuration for their own studies.
@@
Leveraging M3BenchMaker, we create M3Bench, a comprehensive benchmark for task-oriented whole-body motion generation in household environments (see Fig. 1). M3Bench features challenging object rearrangement tasks that require a mobile manipulator to understand its embodiment, environmental contexts, and task objectives from 3D scans to produce coordinated motion for picking or placing objects across diverse scenes. It comprises 30k object rearrangement tasks involving 32 distinct object types across 119 diverse household scenes, covering a wide spectrum of environmental and task objective constraints on robot embodied motion. The benchmark provides expert demonstrations for each task generated by M3BenchMaker, along with additional metadata including natural language task instructions, panoptic maps, and ego-centric camera videos to facilitate research in related areas of embodied AI, such as embodied instruction following etc.
@@
To facilitate in-depth evaluation of motion generation from 3D scans for mobile manipulation, M3Bench incorporates various task splits to assess generalization across different dimensions, such as novel scenes and objects. We leverage a realistic physics simulation platform [14] to evaluate the feasibility of generated motion trajectories and determine whether the robot can physically grasp the object or place it steadily at the desired location. Furthermore, our thorough experiments reveal that both hybrid approaches combining motion planning with affordance prediction and direct imitation learning methods struggle to effectively tackle mobile manipulation tasks across diverse 3D environments. This underscores the necessity of our benchmark for advancing research in whole-body motion generation for mobile manipulation in 3D scenes, paving the way for more adaptive and capable robotic systems in complex environments.
@@
In this work, we make three key contributions as follow:
@@
We introduce, M3Bench, a comprehensive benchmark for task-oriented whole-body motion generation, facilitating the learning and evaluation of mobile manipulation in household environment.
@@
We developed, M3BenchMaker, an automatic data generation tool that procedually generate whole-body motion trajectories from high-level task instructions, allowing users to easily generate data samples with customized robot and scene configurations.
@@
We provided an in-depth evaluation of motion generation from 3D scans for mobile manipulation, revealing weaknesses of current arts in promoting future research in mobile manipulation across diverse 3D scenes.
These contributions collectively advance the state of the art in mobile manipulation, providing the tools and insights necessary to develop more capable and adaptive robotic systems for real-world applications.
II Related Work
Recent benchmarks for Embodied AI have focused on training agents for household activities [26, 27], often simplifying actions to symbolic operations [28] or navigation [5], which limits real-world applicability. While more recent benchmarks allow fixed-base manipulators to interact with objects via realistic simulation [17, 11, 20] or mobile agents to navigate and manipulate in diverse scenes [19, 24, 4], they lack expert data for whole-body motion in 3D scenes. Our M3Bench addresses this gap with a unique collection of whole-body motions for mobile manipulators in various 3D scenes, enabling models to comprehend environmental constraints and generate coordinated motions in complex environments. Table I compares M3Bench with other existing benchmarks.
In addition, unlike existing motion benchmark tools that focus on stationary manipulation [23] in simplified environments, our M3BenchMaker could procedually generate whole-body motion trajectories for mobile manipulators in diverse 3D scenes, facilitating practical model development for real-world tasks.
The proposed M3Bench and tool offer benefits for related fields in embodied AI. In visual affordance, recent studies focus on grasping poses [29, 7, 30] and functional parts [8, 9, 31, 32], but such object-centric modeling often lead to infeasible trajectories in complex scenes. Our whole-body motion data in 3D scenes enables modeling context-aware embodied visual affordance. Similarly, in embodied instruction following, existing benchmarks focus on navigation [5] or symbolic actions [16, 33]. Our M3BenchMaker allows researchers to generate whole-body motion data, facilitating models to produce low-level motion directly from language instructions.
III The M3BenchMaker
@@
Diverse whole-body motion trajectories for mobile manipulators in complex 3D environments is crucial for advancing embodied AI. However, collecting expert demonstrations for training models are usually time-consuming and challenging. To address this, we introduce M3BenchMaker, a user-friendly tool that streamlines the generation of whole-body motion trajectories in 3D scenes, significantly reducing the time and effort required to create large-scale datasets for mobile manipulation tasks in various environments. Fig. 3 illustrates the architecture of the M3BenchMaker.
III-A Task Builder
@@
The task builder serves as the primary user interface, allowing users to define manipulation tasks using high-level action commands such as picking, placing, and reaching. Users no longer need to manually specify grasping poses, placement locations, base positions, or create optimization programs for motion trajectories. To define a task, users simply select target object links from the scene URDF, set the robot’s initial position, and specify the desired action types. The task builder then creates an instance of the data generation pipeline, integrating subsequent modules to procedually generate whole-body motion trajectories. For enhanced dataset diversity, the task builder supports data augmentation via the Conditional Scene Sampler (see Sec. III-B). This feature facilitates the training and evaluation of embodied AI models in complex environments by generating varied scenarios from a single task definition.
III-B Conditional Scene Sampler
@@
where is a set of supporting plane candidates, denotes polygon area, computes intersection, and projects bottom surface points onto the supporting plane, and are distance and angle thresholds. Eq. 1 defines the contact ratio, while Eqs. 2 and 3 enforce alignment and distance constraints. We utilize the method in [34] to extract surface planes and solve the optimization problem by iteratively identifying the plane that maximizes Eq. 1 while satisfying the constraints.
III-C Goal Configuration Generator
@@
Detailed in Alg. 1, our algorithm iteratively selects and updates the sampling probability of candidates based on their feasibility scores. It utilizes a K-D tree for efficient neighbor searching and initializes feasibility scores using the candidates’ energy values. By concentrating sampling in promising regions of the goal configuration space while maintaining exploration, the algorithm significantly reduces the number of expensive feasibility checks required to identify viable goal configurations.
III-D VKC Problem Generator
@@ IV-A Simulation Environment
IV-B Task Design and Variations
@@
The task pool encompasses a wide range of mobile manipulation scenarios, featuring 32 object types with varying properties across 119 diverse household scenes. Each scene presents unique layouts, furniture arrangements, and obstacle configurations. Tasks are generated by selecting appropriate objects and placement locations based on scene categories. We employ the conditional scene sampler (Sec. III-B) to generate various initial configurations, further challenging the robot to generate coordinated whole-body motions while adapting to environmental constraints and task objectives.
IV-C Data Collection
@@ IV-D Benchmark
V Experiments
V-A Experimental Setup
Models for M3Bench. Due to the lack of existing models for whole-body motion generation in mobile manipulation within 3D scenes, we adapt three state-of-the-art approaches to our benchmark:
Implementation Details. For MNet and MFormer, we generate 3D scans from scene URDF. To enhance learning tractability, we apply a perception bounding box around the robot and target object to crop the scans, focusing the model’s attention on relevant spatial information. We train MNet and MFormer on the Train split and perform model selection on Val. In constrast, as ModularMP does not involve learning procedure, we evaluate it directly on the Test and Novel splits. To simplify the optimization problem in ModularMP, we ignore collisions between the end-effector and target object during motion planning, as considering these collisions would frequently result in infeasible trajectories.
V-B Experimental Results
@@
The experimental results are summarized in Tab. IV. Trajectories are evaluated in Isaac Sim using metrics described in Tab. III. Particularly, for the ModularMP model, when motion planning fails to solve the problem (i.e., optimization does not converge), we consider it as a failure instance.
@@
Across Models. Comparing the baseline models, we observed that ModularMP outperforms other baselines in both pick and place tasks, demonstrating higher success rates and lower distances to goal. This aligns with our hypothesis that integrating conventional motion planning with affordance prediction would better generalize across diverse 3D scenes. However, ModularMP’s superior performance comes at the cost of significantly increased computation time, primarily due to optimization complexity in large-scale 3D environments. Additionally, its effectiveness is constrained by the quality of predicted grasp and placement poses, as inappropriate predictions may lead to optimization failures or environmental collisions (see Fig. 2(a)). Despite ModularMP’s relative success, the overall low success rates indicate that combining conventional motion planning with affordance prediction is insufficient.
On the other hand, learning-based models MNet and MFormer, while more time-efficient, struggle to produce feasible whole-body trajectories in complex scenes. Their performance is particularly poor, with near-zero success rates even in the Base split containing familiar objects and scenes. The generated trajectories often violate joint limitations and cause collisions, indicating a fundamental inability to adapt to complex 3D environments and produce feasible whole-body motions. These findings highlight the persistent challenge of generating whole-body motion trajectories for robots in complex 3D scenes. Further research is needed to enhance robot capabilities in manipulating complex environments.
@@
Across Tasks. The experiment results reveal distinct performance patterns between pick and place tasks. While ModularMP maintains better performance in both tasks, its success rates significantly drop in place tasks, and all models require more time to generate trajectories for the place tasks. This discrepancy suggests that generating coordinated whole-body motion trajectories for placing objects is more challenging than for picking objects, as it involves additional constraints such as stable placement locations, appropriate object orientation, and reachable motion trajectory. The increased complexity of place tasks explains the lower success rates and longer execution times observed across all models.
@@
On Generalization. While the learning-based methods MNet and MFormer exhibit near-zero success rates across all test splits, including the Base and generalization splits, we observe significantly higher distances from the end-effector to goal in the Novel splits, indicating these models’ struggle to adapt to unseen objects and scenes. Particularly, the distance to goal in unfamiliar scenes (Novel Scene and Novel Scenario splits) exceeds that of the Novel Object split, suggesting that the impact of novel scenes is more significant than novel objects. In contrast, the conventional motion planning method ModularMP demonstrates consistent performance across all splits. Nevertheless, their low success rates across splits underscore the considerable difficulties of mobile manipulation in diverse household environments.
Remarks. Our experimental results reveal two crucial insights:
@@
While combining motion planning with affordance prediction shows consistent performance across all splits, its overall success rates remain low. This indicates that even advanced hybrid approaches fail to fully address the challenges of mobile manipulation in diverse 3D scenes, underscoring the urgent need for more sophisticated models capable of tackling complex mobile manipulation tasks.
Current arts of learning methods that directly imitate whole-body trajectories fail dramatically when applied to mobile manipulation across large-scale 3D scenes and objects. Their near-zero success rates across all splits highlight the necessity for significant advancements in two key areas: (a) the development of increasingly fine-grained representations for perceptual inputs to better capture the complexity of 3D environments, and (b) more sophisticated modeling of continuous whole-body motions to generate feasible trajectories in challenging scenarios.
VI Conclusion
We introduced M3Bench, a comprehensive benchmark for whole-body motion generation in mobile manipulation tasks across diverse 3D environments, featuring 30k object rearrangement tasks in 566 household scenes. We developed M3BenchMakerto efficiently generate whole-body motion trajectories from high-level instructions. Our experiments revealed significant limitations in current approaches, with even the best-performing hybrid method achieving low success rates and learning-based methods struggling dramatically. While M3Benchprovides a solid foundation for future research, several challenges remain: bridging the gap to real-world scenarios, extending to long-horizon tasks, expanding language instruction diversity, developing more sophisticated models for 3D context understanding and coordinated motion generation, facilitating research into versatile agents, and scaling up demonstrations. Addressing these challenges will be crucial for advancing the field towards more capable and adaptive robotic systems in diverse, complex, unstructured environments.

### v3

@@ Abstract
We propose M3Bench, a new benchmark for whole-body motion generation in mobile manipulation tasks. Given a 3D scene context, M3Bench requires an embodied agent to reason about its configuration, environmental constraints, and task objectives to generate coordinated whole-body motion trajectories for object rearrangement. M3Bench features 30,000 object rearrangement tasks across 119 diverse scenes, providing expert demonstrations generated by our newly developed M3BenchMaker, an automatic data generation tool that produces whole-body motion trajectories from high-level task instructions using only basic scene and robot information. Our benchmark includes various task splits to evaluate generalization across different dimensions and leverages realistic physics simulation for trajectory assessment. Extensive evaluation analysis reveals that state-of-the-art models struggle with coordinating base-arm motion while adhering to environmental and task-specific constraints, underscoring the need for new models to bridge this gap. By releasing M3Bench and M3BenchMaker at https://zeyuzhang.com/papers/m3bench, we aim to advance robotics research toward more adaptive and capable mobile manipulation in diverse, real-world environments.
Index Terms:
Mobile Manipulation, Whole-body Motion Generation, Embodied AI
@@ I Introduction
Humans possess an innate ability to manipulate their environment with remarkable flexibility and coordination, seamlessly integrating locomotion and manipulation. In contrast, robots still struggle to achieve this level of adaptability and proficiency in mobile manipulation. Current learning-based models and motion planning methods for mobile manipulators often address individual subproblems in isolation, such as navigating to waypoints, manipulating with a fixed mobile base, or grasping objects. However, neglecting the potential of coordinated whole-body motion can lead to misalignment between module outputs and task constraints. For instance, in a typical object-fetching task, a navigable position near the target object may still be impossible for the arm to reach the object, or a feasible grasp pose may become unachievable due to collisions with surrounding objects (see Fig. 2(a)). These limitations underscore the necessity of coordinating whole-body motion with a comprehensive understanding of robot embodiment, environmental context, and task objectives to enable effective mobile manipulation in complex 3D scenes.
To generate whole-body motion for mobile manipulation tasks, there is an ongoing debate regarding the effectiveness and limitations of model-based motion planning methods versus data-driven learning-based models. While motion planning can produce complex whole-body mobile manipulation skills [1], its effectiveness and generalizability in real-world scenarios are constrained by its reliance on perfect environmental knowledge [2, 3, 4, 5] or predefined goal configurations (e.g., grasp poses [6, 7, 8]). On the other hand, learning-based models have shown promising results in execution under perceptual and action noise, chaining primitive skills, and adapting to certain environmental variations [9]. However, they have yet to demonstrate robust base-arm coordination with situated goal configurations (e.g., achieving specific grasps or object placements). Learning complex mobile manipulation tasks requires datasets that capture whole-body motions in 3D scenes, yet such datasets remain scarce due to the challenges in generating whole-body motion data. Furthermore, evaluating learned models necessitates a standardized environment for fair benchmarking.
Table I presents recent state-of-the-art benchmarks for Embodied AI and robotics. Many of these benchmarks simplify actions to symbolic operations [10, 11, 12] or navigation [13], lacking physical interaction with the environment. While more recent benchmarks enable fixed-base manipulators to interact with objects in realistic simulations [14, 15, 16, 17] or allow mobile agents to navigate and manipulate in 3D scenes [18, 19, 20, 21], they often overlook the necessity of coordinating base and arm motions.
@@
To address the need for generating whole-body motions in mobile manipulation, we introduce M3Bench, a comprehensive benchmark that features challenging object rearrangement tasks that require a mobile manipulator to reason about its embodiment, environmental context, and task objectives to generate coordinated motions for picking and placing objects in diverse household scenes (see Fig. 1). M3Bench comprises 30,000 object rearrangement tasks involving 32 distinct object types across 119 household scenes, covering a broad spectrum of task objectives and environmental constraints relevant to embodied mobile manipulation. Additionally, it includes rich metadata, such as natural language task instructions, panoptic maps, and egocentric camera videos, making it a valuable resource for related research in Embodied AI, such as embodied instruction following and human-AI collaboration.
Leveraging M3Bench, we developed M3BenchMaker (see Fig. 2(b)), an automatic data generation tool designed to produce whole-body motion trajectories as expert demonstrations for robot learning. M3BenchMaker procedurally generates coordinated trajectories from high-level task instructions, requiring only the action type, object link, and the Unified Robot Description Format (URDF) of the scene and robot. It employs an energy-based model to predict grasp pose and placement candidates [22], and it leverages an advanced virtual kinematics technique [1] to compute coordinated whole-body motion trajectories (see Sec. II for details). This tool not only addresses the scarcity of high-quality whole-body mobile manipulation data but also allows researchers to generate additional samples customized to specific robot and scene configurations for their own studies.
@@
To enable in-depth evaluation of motion generation from 3D scans for mobile manipulation, M3Bench incorporates various task splits to assess generalization across different dimensions, such as novel scenes and objects. We utilize a realistic physics simulation platform [23] to evaluate the feasibility of generated motion trajectories, ensuring that the robot can physically grasp objects and place them stably at the desired locations. Furthermore, our benchmarking reveals that sampling- and optimization-based motion planning methods [24, 25], even when augmented with affordance prediction, as well as learning-based autoregressive planning and generative AI techniques [9, 26, 27], struggle to effectively solve mobile manipulation tasks when required to account for goals such as grasp poses for picking and placement locations for placing actions. After integrating action goals into motion generation, learning-based methods outperform modularized motion planning in computational efficiency and simplicity of problem setup but still lag in motion accuracy. This underscores the importance of high-quality whole-body mobile manipulation data generated by tools like M3BenchMaker and highlights the necessity of M3Bench for advancing research in whole-body motion generation for mobile manipulation in 3D scenes.
@@
Contribution
We make the following contributions:
@@
We introduce M3Bench for benchmarking task-oriented whole-body motion generation for mobile manipulation in household environment, and we provide assets required for testing traditional planning-based methods or learning-based methods.
@@
We develop M3BenchMaker, an automatic whole-body motion generation tool based on high-level task instructions, which can be easily customized for different robot and scene configurations.
@@
We provide an in-depth evaluation of motion generation from 3D scans for mobile manipulation, revealing weaknesses of current arts in promoting future research in mobile manipulation across diverse 3D scenes.
Overview
The remainder of this paper is organized as follows. Sec. II describes the key components of M3BenchMaker. Sec. III details the development of the environment and the benchmarking setup. We implement multiple methods for mobile manipulation and discuss their performance in Sec. IV, and we conclude the paper in Sec. V.
II The M3BenchMaker
@@
Diverse whole-body motion trajectories for mobile manipulators in complex 3D environments is crucial for advancing embodied AI. However, collecting expert demonstrations for training models are usually time-consuming and challenging. To address this, we introduce M3BenchMaker, a user-friendly tool that streamlines the generation of whole-body motion trajectories in 3D scenes, significantly reducing the time and effort required to create large-scale datasets for mobile manipulation tasks in various environments. Notably, M3BenchMaker is adaptable to different robot and scene given the URDF files that describe the configurations, enabling researchers to generate customized whole-body motion trajectories for their specific research needs. Fig. 3 illustrates the architecture of the M3BenchMaker.
II-A Task Builder
@@
The task builder serves as the primary user interface, allowing users to define manipulation tasks using high-level action commands such as picking, placing, and reaching. Users no longer need to manually specify grasping poses, placement locations, base positions, or create optimization programs for motion trajectories. To define a task, users simply select target object links from the scene URDF, set the robot’s initial position, and specify the desired action types. The task builder then creates an instance of the data generation pipeline, integrating subsequent modules to procedually generate whole-body motion trajectories. For enhanced data diversity, the task builder supports data augmentation via the Conditional Scene Sampler (see Sec. II-B). This feature facilitates the training and evaluation of embodied AI models in complex environments by generating varied scenarios from a single task definition.
II-B Conditional Scene Sampler
@@
where is a set of supporting plane candidates, denotes polygon area, computes intersection, and projects bottom surface points onto the supporting plane, and are distance and angle thresholds. Eq. 1 defines the contact ratio, while Eqs. 2 and 3 enforce alignment and distance constraints. The complete sampling procedure is detailed in Alg. 1. We utilize the method in [3] to extract surface planes and solve the optimization problem by iteratively identifying the plane that maximizes Eq. 1 while satisfying the constraints.
II-C Goal Configuration Generator
@@
Detailed in Alg. 2, our algorithm iteratively selects and updates the sampling probability of candidates based on their feasibility scores. It utilizes a K-D tree for efficient neighbor search and initializes feasibility scores using the candidates’ energy values. When a candidate fails the feasibility check, the feasibility scor es of its neighbors, identified via the K-D tree within a specified distance, are halved during the update. By concentrating sampling in promising regions of the goal configuration space while maintaining exploration, the algorithm significantly reduces the number of expensive feasibility checks required to identify viable goal configurations.
II-D VKC Problem Generator
@@ III-A Simulation Environment
III-B Task Design and Variations
@@
The task pool encompasses a wide range of mobile manipulation scenarios, featuring 32 object types with varying properties across 119 diverse household scenes. Each scene presents unique layouts, furniture arrangements, and obstacle configurations. Tasks are generated by selecting appropriate objects and placement locations based on scene categories. We employ the conditional scene sampler (Sec. II-B) to generate various initial configurations, further challenging the robot to generate coordinated whole-body motions while adapting to environmental constraints and task objectives.
III-C Data Collection
@@ III-D Benchmark
IV Experiments
IV-A Experimental Setup
Models for M3Bench. Due to the lack of existing models for whole-body motion generation in mobile manipulation within 3D scenes, we adapt five state-of-the-art approaches to our benchmark:
ModMP[S]: Similar to ModMP[O], we replace the planer with a sampling-based planner RRT-Connect [37].
Implementation Details. For MNet, MDiffusion and MFormer, we generate 3D scans from scene URDF. To enhance learning tractability, we apply a perception bounding box around the robot and target object to crop the scans, focusing the model’s attention on relevant spatial information. We train MNet, MDiffusion and MFormer on the Train split and perform model selection on Val. In constrast, as ModMP[O]and ModMP[S] does not involve learning procedure, we evaluate it directly on the Test and Novel splits. To simplify the optimization problem in ModMP[O] and ModMP[S], we ignore collisions between the end-effector and target object during motion planning, as considering these collisions would frequently result in infeasible trajectories.
IV-B Experimental Results
@@
The experimental results are summarized in Tab. IV. Trajectories are evaluated in Isaac Sim using metrics described in Tab. III. Particularly, for the ModMP[O] and ModMP[S] model, when motion planning fails to solve the problem (i.e., optimization does not converge), we consider it as a failure instance.
@@
Across Models. ModMP[O] consistently outperforms other approaches in most pick-and-place tasks, achieving higher success rates and closer goal distances. This confirms our hypothesis that integrating conventional motion planning with affordance prediction generalizes effectively across diverse 3D scenes. While the sampling-based ModMP[S] performs comparably to the optimization-based ModMP[O], it requires substantially more processing time. However, ModMP[O]’s superior performance comes with increased computation time due to optimization complexity in large-scale environments. Its effectiveness depends heavily on grasp and placement pose prediction quality; inaccuracies can lead to optimization failures or collisions (see Fig. 2(a)). Despite promising results, the overall low success rates indicate that this integration alone is insufficient for robust performance.
In contrast, learning-based models are more time-efficient but often fail to generate feasible solutions in unseen scenarios. Specifically, while MDiffusion achieves comparable performance to planning-based methods, its success drops significantly in the Novel splits. Unlike planning-based methods with hard constraints, learning-based approaches frequently produce trajectories violating joint limitations and prone to collisions. MNet and MFormer achieve minimal success in test splits, suggesting that direct adaptation of stationary manipulation models to mobile manipulation is infeasible without sufficient model capacity to capture environmental complexity. These findings underscore the challenge of generating whole-body motion trajectories in complex 3D environments and highlight the need for more sophisticated mobile manipulation models.
@@
Across Tasks. The experiment results reveal distinct performance patterns between pick and place tasks. While ModMP[O] maintains better performance in both tasks, its success rates significantly drop in place tasks, and all models require more time to generate trajectories for the place tasks. This discrepancy suggests that generating coordinated whole-body motion trajectories for placing objects is more challenging than for picking objects, as it involves additional constraints such as stable placement locations, appropriate object orientation, and reachable motion trajectory. The increased complexity of place tasks explains the lower success rates and longer execution times observed across all models.
@@
On Generalization. While MDiffusion achieves comparable performance to planning-based methods in the Base split, its performance deteriorates in the Novel splits,indicating persistent generalization challenges for learning-based models. Particularly, the distance to goal in unfamiliar scenes (Novel Scene and Novel Scenario splits) exceeds that of the Novel Object split, suggesting that the impact of novel scenes is more significant than novel objects. The conventional planning-based method, in contrast, maintains consistent performance across all splits, though its relatively low success rates across the board underscore the inherent complexity of mobile manipulation tasks in diverse household environments.
Remarks. Our experiments reveal two key insights:
@@
Although combining motion planning with affordance prediction demonstrates consistent performance across all splits, its overall success rates remain low. This highlights the limitations of hybrid approaches and the need for holistic solutions to mobile manipulation in 3D scenes.
Mobile manipulation requires models with greater expressiveness and generalization than stationary tasks. Poor performance in unseen scenarios (i.e., Novel splits) underscores the need for (a) fine-grained perceptual representations and (b) advanced whole-body motion models for feasible trajectories.
V Conclusion
We introduced M3Bench, a comprehensive benchmark for whole-body motion generation in mobile manipulation tasks across diverse 3D environments, featuring 30k object rearrangement tasks in 119 household scenes. M3Bench provides a standardized platform for both planning and learning communities. Through comprehensive evaluations of state-of-the-art models, we highlighted the persistent challenges in generating coordinated base-arm motion trajectories that satisfy both environmental constraints and task objectives. Furthermore, we developed M3BenchMaker, a tool designed to efficiently generate whole-body motion trajectories from high-level instructions, which can serve as a valuable resource for researchers in their own studies. We hope M3Bench opens new opportunities for robotics research and catalyzes progress toward developing more adaptive and capable embodied agents.